OpenStudy (anonymous):

which equation is the equation of a line that passes through (-10 3) and is perpendicular to y=5x-7

OpenStudy (anonymous):

Well, for a line to be perpendicular, it just needs to have it's slope be the negative inverse. That is to say \(-1/5\) will be the slope in this case.

OpenStudy (anonymous):

So we end up with: \[ \frac{y-(3)}{x-(-10)} = -\frac 15 \]And we just have to solve for \(y\) if we want it in slope point formula.

OpenStudy (anonymous):

Refer to the attached plot.

OpenStudy (anonymous):

Thank you! :D